OpenStudy (anonymous):

I really need help on trig substitution I get it but when I try to simplify its really hard for me.

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ \sqrt{x^2+4} }\]

OpenStudy (anonymous):

can you please show all your steps of simplifying

OpenStudy (anonymous):

I managed to simplify it down to: \[\int\limits_{}^{}\sec \theta\] where theta equals:\[\arctan(\frac{ x }{ 2 })\]

OpenStudy (anonymous):

but the integral of sec theta is really long and becomes HUGE when you back-substitute the xs

OpenStudy (anonymous):

i simplified it down as far as possible

OpenStudy (anonymous):

What you've done is correct, well I get the same simplification

OpenStudy (anonymous):

but then do i substitute arctan(x/2) for theta?

OpenStudy (anonymous):

after doing the integral

OpenStudy (anonymous):

Then using substitution I would let: u=arctan(x/2) then you have: \[\int\limits_{}^{}\sec(u)du\] multiply top and bottom by tan(u)+sec(u) \[=\int\limits_{}^{}\frac{ \sec^2(u)+\tan(u)\sec(u) }{ \tan(u)+\sec(u) }du\] let p=tan(u)+sec(u) then dp=(sec^2(u))+tan(u)sec(u) du \[=\int\limits_{}^{}\frac{ 1 }{ p }dp = log(p) + constant\] then sub back in that p=tan(u)+sec(u) and sub your 'u' back in too

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

yeah sorry, the final solution should be in terms of 'x'

OpenStudy (anonymous):

i also know that \[\int\limits_{}^{}\sec \theta=-\log(\cos(θ/2)-\sin(θ/2))+\log(\cos(θ/2)+\sin(θ/2))\]

OpenStudy (anonymous):

so now i would sub back in arctan(x/2) for theta right?

OpenStudy (anonymous):

and then simplify

OpenStudy (anonymous):

i got: \[1/4 \sqrt{(x^2+4)} x+\sinh^{-1})(\frac{ x }{ 2 })+c\]

OpenStudy (anonymous):

something like that

OpenStudy (anonymous):

ok thx c ya

OpenStudy (anonymous):

the final solution I got was:\[\log(\frac{ 1 }{ 2 }(\sqrt{x^2+4}+x) + constant\]